Problem 1:

Consider a

$

1000

bond with a coupon rate of 10% and annual coupons. The

par value is $1,000, and the bond has 5 years to maturity. The yield to maturity is

9

%.

For

each question,

s

how your

wo

rk/

calculations

.

A.

What is the

present value of coupons

?

B.

What is the

present value

of face value (i.e. par value)

?

C.

What is the value of the bond

?

D.

Is it a prem

i

um

or

discount

bond?

Problem

2

:

Consider a

$

1000

bond with a coupon rate of

7.5

% and annual coupons. The

par value is $1,000, and the bond has

10

years to maturity. The yield to maturity is

10

%.

For

each question,

s

how your

wo

rk/

calculations

.

A.

What is the

present value of coupons

?

B.

What is the

present value

of face value (i.e. par value)

?

C.

What is the value of the bond

?

D.

Is it a prem

i

um

or

discount

bond?

Grading Criteria:

Your submission

will be evaluated based on the following criteria (1%):

Content: adequate coverage of topic, logic of arguments,

relevance and indication

of all used resources.

Presentation/Clarity: use of appropriate headings, appropriate organization, and

ease of understanding.

Accuracy: free of spelling, grammar, and content errors.
Answer #1

Problem 1:

Consider a $1000 bond with a coupon rate of 10% and annual coupons. The par value is $1,000, and the bond has 5 years to maturity. The yield to maturity is 9%.

For each question, show your work/calculations

A. Present value of coupons is calculated below:

We are given the following information

PMT $100.00 = 10% x 1000
r 9.00%
n 5

We need to solve the following equation to arrive at the required PV
PV = PMT x 1-(1+r)-n 1-(1+0.09) -5 PV = 100 x 0.09 PV = 388.97

So the PV is $388.97

B. Present value of face value is calculated below:

We are given the following information:

r 9.00%
n 5
FV $              1,000.00

We need to solve the following equation to arrive at the required FV
FV = PV X (1+r) 1000 = PV X (1+0.09) PV = 649.93

So the PV is 649.93

C. Value of the bond = sum of the PV of the coupons + sum of the PV of par value

Value of the bond = 388.97 + 649.93

Value of the bond = 1038.9

We can even create a schedule for the PV as follows:

Year CF Discount Factor Discounted CF
1 $    100.00 1/(1+0.09)^1= 0.917431193 0.91743119266055*100= $       91.74
2 $    100.00 1/(1+0.09)^2= 0.841679993 0.84167999326656*100= $       84.17
3 $    100.00 1/(1+0.09)^3= 0.77218348 0.772183480061064*100= $       77.22
4 $    100.00 1/(1+0.09)^4= 0.708425211 0.708425211065196*100= $       70.84
5 $1,100.00 1/(1+0.09)^5= 0.649931386 0.649931386298345*1100= $     714.92
PV = Sum of all Discounted CF $ 1,038.90

D.Is it a premium or discount bond?

As the PV of the bond is greater than the par value, so this is a premium bond. This can even be arreived at by the fact that the discount rate < coupon rate, when this is the case, then the bond is a premium bond

Problem 2

Consider a $1000 bond with a coupon rate of 7.5% and annual coupons. The par value is $1,000, and the bond has

10 years to maturity. The yield to maturity is 10%. For each question, show your work/ calculations

A. Present value of coupons

We are given the following information

PMT $75.00 = 0.075 x 1000
r 10.00%
n 10

We need to solve the following equation to arrive at the required PV
PV = PMTX 1-(1+r)-n PV = 75 x 1-(1 + 0.1) -10 0.1 PV = 460.84

So the PV is $460.84

B. Present value of face value:

We are given the following information:

r 10.00%
n 10
FV $              1,000.00

We need to solve the following equation to arrive at the required PV
FV = PV X (1+r) 1000 = PV x (1 +0.1)10 PV = 385.54

So the PV is 385.54

C. The value of the bond = 385.54 + 460.84 =  $846.39

Year CF Discount Factor Discounted CF
1 $      75.00 1/(1+0.1)^1= 0.909090909 0.909090909090909*75= $       68.18
2 $      75.00 1/(1+0.1)^2= 0.826446281 0.826446280991735*75= $       61.98
3 $      75.00 1/(1+0.1)^3= 0.751314801 0.751314800901578*75= $       56.35
4 $      75.00 1/(1+0.1)^4= 0.683013455 0.683013455365071*75= $       51.23
5 $      75.00 1/(1+0.1)^5= 0.620921323 0.620921323059155*75= $       46.57
6 $      75.00 1/(1+0.1)^6= 0.56447393 0.564473930053777*75= $       42.34
7 $      75.00 1/(1+0.1)^7= 0.513158118 0.513158118230706*75= $       38.49
8 $      75.00 1/(1+0.1)^8= 0.46650738 0.466507380209733*75= $       34.99
9 $      75.00 1/(1+0.1)^9= 0.424097618 0.424097618372485*75= $       31.81
10 $1,075.00 1/(1+0.1)^10= 0.385543289 0.385543289429531*1075= $     414.46
NPV = Sum of all Discounted CF $     846.39

D. As the PV of the bond is lower than the par value, so this is a discount bond. This can even be arreived at by the fact that the discount rate > coupon rate, when this is the case, then the bond is a discount bond